Derive a formula for
n

i= i² a telescoping sum with terms f(I) = i³


n
∑ (3i-1)²
i=1


n
∑ (2i + 1)²
i=1
1
Expert's answer
2020-12-03T07:46:13-0500

"\sum_{i=1}^{n} i"

= 1+2+3+......+n

= n(n+1)/2 by formula of sum of AP


(i+1)³ - i³ = 3i²+3i+1

"\sum_{i=1}^{n} (i+1)^3 -i^3 = \\\sum_{i=1}^{n} 3i²+\sum_{i=1}^{n} 3i + \sum_{i=1}^{n} 1"

=>

"\sum_{i=1}^{n} (i+1)^3 -i^3 =\\ 3\sum_{i=1}^{n} i²+3\sum_{i=1}^{n} i + \sum_{i=1}^{n} 1"

=>

2³-1³+3³-2³+4³-3³....(n+1)³-n³ =

"3\sum_{i=1}^{n} i²+3\sum_{i=1}^{n} i + \sum_{i=1}^{n} 1"

=>

(i+1)³ - i³ = "3\sum_{i=1}^{n} i²+3\sum_{i=1}^{n} i + \sum_{i=1}^{n} 1"

So

"3\sum_{i=1}^{n} i² \\= (n+1)³ - 1 - 3\sum_{i=1}^{n} i - \sum_{i=1}^{n} 1"

= n³ + 3n² + 3n + 1 - 1 - 3n(n+1)/2 -n

= "\frac{2n³+3n²+n}{2}"

= "\frac{n(n+1)(2n+1)}{2}"

So "\sum_{I=1}^{n} i² =" "\frac{n(n+1)(2n+1)}{6}"


Now

(i+1)⁴ - i⁴= 4i³+6i²+4i+1

"\sum_{1}^{n}{(i+1)⁴ - i⁴}= 4\sum_{1}^{n} i³+6\sum_{1}^{n} i²+4\sum _{1}^{n}i+\sum_{1}^{n} 1"

=> 2⁴-1⁴+3⁴-2⁴.....(n+1)⁴-n⁴

= 4"\sum_{1}^{n} i³" + 6"\frac{n(n+1)(2n+1)}{6}" + 4 "\frac{n(n+1)}{2}" +n

=> 4"\sum_{1}^{n} i³"

= (n+1)⁴-1-n(n+1)(2n+1) - 2n² - 2n - n

= (n+1)⁴ - (n+1)(2n²+n+2n+1)

= (n+1)⁴ - (n+1)(2n+1)(n+1)

= (n+1)²(n²+2n+1-2n-1)

= n²(n+1)²

So

"\sum_{1}^{n} i³" = "[\frac{n(n+1)}{2}]^2"


"\sum_{i=1}^{n} (3i-1)²" "=\sum_{i=1}^{n} (9i²-6i+1)"

"=9\sum_{i=1}^{n} i²-6\sum_{i=1}^{n} i + \sum_{i=1}^{n} 1"

= 9 "\frac{n(n+1)(2n+1)}{6}" - 6 "\frac{n(n+1)}{2}" + n

= "\frac{n(6n²+3n-1)}{2}"


"\sum_{i=1}^{n} (2i+1)²"

= "\sum_{i=1}^{n} (4i²+4i+1)"

= "4\sum_{i=1}^{n} i²+4\sum_{i=1}^{n} i + \sum_{i=1}^{n} 1"

= 4"\frac{n(n+1)(2n+1)}{6}" + 4"\frac{n(n+1)}{2}" + n

= "\frac{n(4n²+12n+11)}{3}"